OpenStudy (anonymous):

medal && fan ! a food company makes regular and tall soup cans. The area of the base of both cans is 5 cm^2. The volume of the regular can is 40cm^3 The tall can is 2 cm taller. What is the volume of the tall soup can?

OpenStudy (anonymous):

@Jennjuniper

OpenStudy (anonymous):

1st, whats the equation for a cylinder?

OpenStudy (anonymous):

\[\pi r ^{2} h\] or (Base area)(height)

OpenStudy (anonymous):

so. try plugging in the variables you already have.

OpenStudy (anonymous):

@_DimondLife_

OpenStudy (anonymous):

sorry im working on a test right now and its time give me a couple of mins please,

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

yesterday i got a little busy im so sorry , do you have time to help me now ?

OpenStudy (anonymous):

yeah. sure.

OpenStudy (anonymous):

okay so what do i do now ?

OpenStudy (anonymous):

\[\pi r ^{2} h\] take this equation and plug in the variables you already know.

OpenStudy (anonymous):

so its 5 times 10 ?

OpenStudy (anonymous):

plug in the variables to the equation so i can see first.

OpenStudy (anonymous):

i meant to put 40 not 10 sorry

OpenStudy (anonymous):

the \[\pi r ^{2}\] is the area of the base btw.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

because the base area of both cans is 5cm^2

OpenStudy (anonymous):

okay. so the equation is\[40=5h\] right?

OpenStudy (anonymous):

yes , and then i do what ?

OpenStudy (anonymous):

what do you think? remember, we're trying to isolate the variable here.

OpenStudy (anonymous):

do i multiply 40 by 5 or divide it

OpenStudy (anonymous):

whatever the action is between the 5 and the h, you do the opposite.

OpenStudy (anonymous):

40 divided by 5 = 8

OpenStudy (anonymous):

yes! so what is the height of the regular can?

OpenStudy (anonymous):

8 ?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

so. what does it say about the regular can versus the large can?

OpenStudy (anonymous):

it says the tall can is 2 cm taller what is the volume of the tall soup can ?

OpenStudy (anonymous):

well, you need height to find volume. the height is 10 (you did 8+2=10). so what would your equation be? remember, they both have the same base area

OpenStudy (anonymous):

idk ? i dont get this stuff at all

OpenStudy (anonymous):

\[Volume=Base Area \times height\] insert your variables! c'mon, you can do it!

OpenStudy (anonymous):

10 times 5 ?

OpenStudy (anonymous):

yeah. just the other way around is all.

OpenStudy (anonymous):

So the answer is 50 ?